| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| hgfw2: | HG FW#2 at step 192 | 10 | 5 | B34725.021 | TORON | #021 | 05:40:46 | 192 | 0 | 129 |
| sl: | Lamp not on... sl test terminated F1= 9 F4= 269 | 3 | 1 | B34725.021 | TORON | #021 | 13:47:54 | 1 | -1 | |
| hp: | Spectrometer alignment trouble. | 2 | 1 | B34725.021 | TORON | #021 | 15:18:06 | 1 | -1 | |
| hg: | Lamp not detected f1= 5 f4= 4 | 1 | 1 | B34725.021 | TORON | #021 | 18:23:28 | 1 | -1 | |
| re: | Previous RE didn't seem to help, so not doing reset again. Operator intervention required. Quitting to menu. | 1 | 1 | B34725.021 | TORON | #021 | 19:26:41 | 1 | -1 |
| Serial | YYYYJJJ | HP large | max correction |
|---|---|---|---|
| 021 | 2025347 | 11 | 30.0 |
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| ds: | Suppressed oscillation of filter wheel. | 4 | 5 | B34725.021 | TORON | #021 | 15:31:38 | 1 | -1 | |
| pd: | Going to menu | 1 | 99 | B34725.021 | TORON | #021 | 19:26:43 | 1 | -1 |
| _jday | _loc | _serial | _skc | aode | ds | dt | fz | hg | hp | rs | sl | uv | zs |
|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
| 347 | TORONTO | #021 | calo3em,Waiting until ZA = -93 now -101.5,Waiting until ZA = -93 now -103.2,Waiting until ZA = -93 now -105,Waiting until ZA = -93 now -93.3,Waiting until ZA = -93 now -94.9,Waiting until ZA = -93 now -96.5,Waiting until ZA = -93 now -98.2,Waiting until ZA = -93 now -99.9 | 51 | 54 | 2 | 2 | 9 | 21 | 2 | 2 | 4 | 12 |
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
|---|---|---|---|---|
| 021 | 5.4 | 0.6 |